A shaft is made of a steel alloy having an allowable shear stress of ?(allow) = 12 ?si. If the diameter of the shaft is 1.5 in.,

(a) determine the maximum torque T that can be transmitted.

(b) What would be the maximum torque T’ if a 1-in.-diameter hole is bored through the shaft?

(c) Sketch the shear-stress distribution along a radial line in each case. 2. A shaft is made of a steel alloy having an allowable shear stress of Tallow = 12 ksi. If the diameter of the shaft is 1.5
Answer #1

a) 16T Tallow TO 3 Tallowa 12 ksi CGiven) 16xT 3 12X10 = TX(1.5) Dolor 3 (1.53 X TX 12x10 TE 16 16-m T = 795201 56 lbin Toaqu

7 do 3 (164) * Tallow T = 16. *) x 12X103 5 ax (165)*x (6-( %) (6 Ib-io 63 81. 36 T = thant see we can c) Form Torsional equa